OpenStudy (anonymous):

The probability of a person over the age of 50 developing cancer is .06 and the probability that she/he develops heart problems is .09. If a patient over 50 is selected at random, what is the probability that the patient will develop at least one of the diseases? Round your answer to two decimal places.

OpenStudy (anonymous):

help

OpenStudy (zarkon):

P(at least one)=1-P(gets none)=1-P(no C and no H) =1-P(no C)*P(no H) =1-(1-.06)(1-.09)=.1446

OpenStudy (anonymous):

This can't be solved without knowing something about the probability that they will get both. The two probabilities are not independent.

OpenStudy (zarkon):

I believe that is an assumption to be made (that they are independent) based on previous problems asked by hscheide

OpenStudy (anonymous):

Ah, understood.